# Navigating PESQ: Up-to-Date Versions and Open Implementations

### v1

@@ Abstract
Perceptual Evaluation of Speech Quality (PESQ) is an objective quality measure that remains widely used despite its withdrawal by the International Telecommunication Union (ITU). PESQ has evolved over two decades, with multiple versions and publicly available implementations emerging during this time. The numerous versions and their updates can be overwhelming, especially for new PESQ users. This work provides practical guidance on the different versions and implementations of PESQ. We show that differences can be significant, especially between PESQ versions. We stress the importance of specifying the exact version and implementation that is used to compute PESQ, and possibly to detail how multi-channel signals are handled. These practices would facilitate the interpretation of results and allow comparisons of PESQ scores between different studies. We also provide a repository that implements the latest corrections to PESQ, i.e., Corrigendum 2, which is not implemented by any other openly available distribution: https://github.com/audiolabs/PESQ.
@@ 1.1 PESQ history, importance, and caveats
PESQ [1, 2, 3, 4, 5] was developed to predict the mean opinion score (MOS) of a transmitted speech signal by comparing it with a clean reference. PESQ has a long history as an International Telecommunications Union – Telecommunications Sector (ITU-T) recommendation P.862, with its first version published in 2001 and the last document dated 2018, as summarized in Table 1. In 2018, PESQ was withdrawn by the ITU to be replaced by Perceptual Objective Listening Quality Analysis (POLQA) [6, 7]. Despite its withdrawal, PESQ remains a widely used metric for researchers and practitioners. This is evident by conducting a Google Scholar search for the term PESQ, which returns over 4,600 search results for the year 2024, while only a little more than 100 results are found for the term POLQA. More recently, methods have been proposed to predict PESQ scores non-intrusively (i.e., without a clean reference) using a neural network [8, 9]. The popularity of PESQ is due to several reasons; chief among them is the abundance of publicly available implementations and the robustness that PESQ shows across different application domains [10]. Still, given the evolution of PESQ and its versions, the availability of several implementations can lead to confusion and may render PESQ values reported in different works incomparable.
@@ 1.2 Contributions and structure of this work
This work examines PESQ implementations and their differences within the larger context of ITU recommendations. The PESQ versions and implementations considered in this work are summarized in Table 2. The remainder of this work is structured as follows. Sec. 1.3 details the test audio signals, Sec. 2 describes the different versions of PESQ. Sec. 3 discusses the different publicly available implementations. The application to multi-channel signals is discussed in Sec. 4. Finally, additional remarks (Sec. 5) and conclusions (Sec. 6) complete this work.
1.3 Test audio signals
As test audio signals in this paper, we use the Open Dataset for Audio Quality (ODAQ) [11]. It is a dataset of stereo audio signals representing a heterogeneous range of processing types and quality levels. In the following experiments both items with and without speech are considered, with the exception of the correlation analysis with subjective scores in Sec. 4, where only signals with speech are selected. For the experiments in Sec. 3, stereo signals are passively downmixed (i.e., by averaging the two channels) before computing PESQ. In all cases, the signals are downsampled to $16$ kHz using librosa [12] with its default parameters, i.e., soxr_hq.
@@ 2 PESQ Versions
Implementation PESQ version Narrowband raw PESQ (P.862) 8 kHz / 16 kHz Narrowband MOS-LQO (P.862.1) 8 kHz / 16 kHz Wideband MOS-LQO (P.862.2) 16 kHz without Corrigendum 2 Wideband MOS-LQO (P.862.2) 16 kHz with Corrigendum 2 Accepts stereo inputs ITU reference code (ANSI C) (RMSE = 0.58) (RMSE = 0.57) (Reference) (RMSE = 0.52) (Intended?) ludlows/PESQ (Python) (RMSE = 0.03) audiolabs/torch-pesq (PyTorch) (RMSE = 0.14) vBaiCai/python-pesq (Python) Loizou (MATLAB) audiolabs/PESQ (Python)
@@
Figure 1: Mapping from narrowband raw PESQ (P.862) to narrowband MOS-LQO (P.862.1).
@@
The third version of PESQ is P.862.2 [4]. This considers a larger frequency range (50-7000 Hz, named wideband) and is available only for signals with a sampling frequency of 16 kHz. This version only outputs MOS-LQO scores and no raw PESQ scores. Compared to P.862.1, P.862.2 introduces new input filters and a new output mapping to MOS-LQO. Fig. 2 compares narrowband raw PESQ (P.862) to wideband P.862.2 scores, using a sampling frequency of 16 kHz. Differences can be as large as 1.86 with RMSE $=0.58$ on the considered data.
@@
More than a decade after the first edition of P.862.2, the filter coefficients for wideband assessment were acknowledged to be incorrect in P.862 Corrigendum 2 [5]. The corrigendum describes a systematic under-prediction of subjective scores. The under-prediction of P.862.2 is estimated to be $0.8$ on average, and the corrected filter coefficients are provided. Fig. 3 shows the differences between P.862.2 with and without Corrigendum 2. The differences in absolute terms are significant with an RMSE $=0.52$ and a maximum difference of $1.30$ for the audio signals considered.
Despite the acknowledged errors, P.862.2 without Corrigendum 2 seems to have established itself as the de facto standard in the audio signal processing community. This could be explained by the lack of an openly available implementation that directly implements Corrigendum 2. This work fills this gap by introducing a publicly available implementation, audioalabs/PESQ, as explained in the following.
Figure 3: Wideband PESQ (P.862.2) with and without Corrigendum 2. Although the correlation is strong, differences can be as big as 1.30, and RMSE is 0.52 on the considered audio signals.
@@ 3.2 Python and PyTorch
ludlows/PESQ [^2] https://github.com/ludlows/PESQ wraps around the ITU reference code and provides access to P.862.1 and P.862.2 for Python users. Fig. 4 compares P.862.2 as output by the ITU reference code and by ludlows/PESQ, showing that they are basically identical.
audiolabs/PESQ [^3] https://github.com/audiolabs/PESQ is the repository we created by forking from ludlows/PESQ and adding the corrections detailed in Corrigendum 2. It is released as PyPI package pesqc2. This provides the possibility of easily and reliably computing P.862.2 with Corrigendum 2. The introduction of Corrigendum 2 results in large differences compared to ludlows/PESQ as shown in Fig. 3.
audiolabs/torch-pesq [^4] https://github.com/audiolabs/torch-pesq provides a torch implementation of P.862.2 without Corrigendum 2. As depicted in Fig. 5 and as acknowledged by the authors, some items can deviate significantly from the ITU reference implementation. The average error is rather small with RMSE $=0.14$ on the data considered.
@@ 4.1 Reference implementation
The results shown so far are obtained considering mono signals. Indeed, PESQ is developed for mono inputs, and all considered publicly available implementations accept only mono inputs (or batches thereof), with one important exception: The ITU reference code accepts multi-channel inputs. It is unclear if this is intentional. In the reference code, the reading routines load all input file samples. In practice, when a multi-channel audio file is given as input, a mono signal is created, in which channels are interleaved. As a consequence, the reference implementation processes the interleaved multi-channel signals as a long single-channel signal.
Figure 4: ludlows/PESQ compared with the ITU reference implementation. Differences are neglectable.
Figure 5: audiolabs/torch-pesq compared with the ITU reference implementation. Some items deviate significantly from the reference implementation.
4.2 Comparing strategies for applying PESQ to stereo signals
In the following, the processing of stereo signals by the ITU reference code is referred to as stereo. More intuitive strategies for applying PESQ to stereo signals include:
@@ 4.2 Comparing strategies for applying PESQ to stereo signals
    Mono-downmixing the stereo audio signal before computing PESQ (mono dmx);
@@
mono dmx and avg scores are compared in Fig. 6, while mono dmx is compared with stereo in Fig. 7. To better understand the performance of the different strategies for applying PESQ to stereo signals, we analyze the correlation with subjective scores of perceived audio quality. In all cases, we consider the ITU reference implementation of P.862.2, with and without adding the fixes of Corrigendum 2 (c2). As the ground truth, we use the scores collected from two listening tests involving stereo signals:
@@
    The Dialogue Enhancement (DE) subset of ODAQ, containing speech signals panned to the center and mixed with movie-like music and effects. Speech enhancement methods are applied to boost the level of speech. The considered enhancement methods create significantly different levels of quality, which are evaluated by expert listeners. The remaining non-speech items from ODAQ are not considered, nor are the extra scores from [15].
@@
    As a second dataset, we consider the results from the listening test described in [16] and named Dialogue Separation (DS). This work compares different speech enhancement methods to improve the level of speech in studio-like recordings.
The correlation coefficients between the different stereo realizations of PESQ and the subjective speech quality scores are provided in Table 3. Both Pearson’s linear correlation and Kendall’s rank correlation are considered. Kendall’s coefficients are mapped as explained in [10, 17] to make their values more comparable to Pearson’s coefficients.
PESQ provides sensible scores in all cases, with correlation coefficients $\geq 0.73$. This prevents drawing a definitive conclusion on the overall best strategy. The relative good performance of PESQ on the stereo signals of ODAQ DE was also observed in [15], where a comparison to other state-of-the-art objective metrics showed PESQ among the best-performing ones. It is worth noting that P.862.2 results in similar correlation with and without Corrigendum 2. As also observed in Fig. 3, Corrigendum 2 introduces a significant difference in absolute terms, rather than in correlation.
Figure 6: Evaluating each channel independently with the ITU reference implementation and then averaging the output PESQ scores compared with passively mono-downmixed inputs.
Figure 7: Stereo signals input to the ITU reference implementation compared with passively mono-downmixed inputs.
5 PESQ for Training
Some works use PESQ as part of the training strategy or validation metrics for Deep Neural Network (DNN)-based speech enhancement (SE), e.g., [18, 19]. In this case, extra care should be taken when using PESQ for the performance evaluation of the trained DNN. For example, in [20], a listening test is reported, which compares SE methods, including a generative method optimized on PESQ. When we compare the subjective scores of this listening test and PESQ P.862.2 outputs, we observe much weaker correlation coefficients than the ones reported so far. Pearson’s correlation is $0.28$, and mapped Kendall’s correlation is $0.31$. This can be explained by the fact that PESQ was overfitted (or fooled). Improvements in PESQ scores did not correspond to improvements in perceptual quality. The problem of DNN-based SE systems that provide high PESQ scores but low perceptual quality is also reported in [21]. In [22], a DNN called PESQNet is proposed to non-intrusively estimate PESQ scores, serving as a differentiable perceptual loss function for optimizing a SE model. To avoid overfitting, the method alternates training between the SE model and PESQNet, effectively guiding the SE model to achieve improved perceptual quality measured by higher PESQ scores.
ODAQ DE [11] DS [16]
ITU ref. P.862.2 (stereo) $73$ $66$ $83$ $85$
ITU ref. P.862.2 (mono dmx) $80$ $73$ $83$ $87$
ITU ref. P.862.2 (avg scores) $76$ $68$ $85$ $90$
ITU ref. P.862.2 c2 (stereo) $74$ $75$ $85$ $88$
ITU ref. P.862.2 c2 (mono dmx) $79$ $77$ $81$ $88$
ITU ref. P.862.2 c2 (avg scores) $75$ $71$ $83$ $90$
Table 3: Pearson’s and mapped Kendall’s correlation between subjective scores from listening tests on stereo speech enhancement and different strategies for applying PESQ to stereo signals.
@@ 6 Conclusion
While PESQ is a widely used speech quality metric, the term PESQ encompasses different proposals and versions. These include: narrowband raw PESQ score (P.862), narrowband MOS-LQO (P.862.1), and wideband MOS-LQO (P.862.2), with and without Corrigendum 2. This corrigendum introduced significant differences, but was not included in any of the numerous publicly available implementations of PESQ. Probably due to this lack, the community seemed to ignore the corrigendum in practice, and P.862.2 without Corrigendum 2 was used most frequently. Our audiolabs/PESQ fills this gap and provides P.862.2 with Corrigendum 2, i.e., the latest version of PESQ.
We also provided practical guidance on the different versions and implementations. We observed that the differences between PESQ versions are generally much more significant than those between different implementations of the same version. A challenge emerges when dealing with multi-channel signals. The reference PESQ implementation by the ITU processes stereo signals and outputs plausible scores. Other more intuitive extensions for stereo processing are discussed.
Finally, it is recommended to specify the PESQ version and implementation when reporting PESQ scores, as this ensures clarity and reproducibility across studies.

### v2

@@ Abstract
Perceptual Evaluation of Speech Quality (PESQ) is an objective quality measure that remains widely used despite its withdrawal by the International Telecommunication Union (ITU). PESQ has evolved over two decades, with multiple versions and publicly available implementations emerging during this time. Different versions and their updates can be overwhelming, especially for new PESQ users. This work provides practical guidance on the different versions and implementations of PESQ. We show that differences can be significant, especially between PESQ versions. We stress the importance of specifying the exact version and implementation that is used to compute PESQ, and possibly to detail how multi-channel signals are handled. These practices would facilitate the interpretation of results and allow comparisons of PESQ scores between different studies. We also provide a repository that implements the latest corrections to PESQ, i.e., Corrigendum 2, which is not implemented by any other openly available distribution: https://github.com/audiolabs/PESQ.
@@ 1.1 PESQ history, importance, and caveats
PESQ [1, 2, 3, 4, 5] was developed to predict the mean opinion score (MOS) of a transmitted speech signal by comparing it with a clean reference. PESQ has a long history as an International Telecommunications Union – Telecommunications Sector (ITU-T) recommendation P.862, with its first version published in 2001 and the last document dated 2018, as summarized in Table 1. In 2018, PESQ was withdrawn by the ITU to be replaced by Perceptual Objective Listening Quality Analysis (POLQA) [6, 7]. Despite its withdrawal, PESQ remains a widely used metric for researchers and practitioners. This is evident by conducting a Google Scholar search for the term PESQ, which returns over 4,600 search results for the year 2024, while only a little more than 100 results are found for the term POLQA. PESQ has been used as part of the training strategy or validation metrics for Speech Enhancement (SE) in various works, e.g., [8, 9, 10, 11], and methods have been proposed to predict PESQ scores non-intrusively (i.e., without a clean reference) using a Deep Neural Network (DNN) [12, 13]. The popularity of PESQ is due to several reasons; chief among them is the abundance of publicly available implementations and the robustness that PESQ shows across different application domains, although its correlation with perceived quality is outperformed by other measures [14].
Given the evolution of PESQ and its versions, the availability of several implementations can lead to confusion and may render the PESQ scores reported in different works incomparable. This paper acknowledges the importance of PESQ in the SE community and aims to provide guidance in navigating its numerous versions and implementations. This paper neither endorses the use of PESQ nor aims to improve it. The validity and accuracy of PESQ as a quality metric are beyond the scope of this work.
@@ 1.2 Contributions and structure of this work
This work examines PESQ implementations and their differences within the larger context of ITU recommendations. The PESQ versions and implementations considered in this work are summarized in Table 2. The remainder of this work is structured as follows. Sec. 2 describes the different versions of PESQ. Sec. 3 discusses the different publicly available implementations. Sec. 4 examines the differences between the versions and the implementations, including a discussion of the used test audio signals. The application to multi-channel signals is discussed in Sec. 5. Section 6 concludes this work.
@@ 2 PESQ Versions
Implementation PESQ version Narrowband raw PESQ (P.862) 8 kHz / 16 kHz Narrowband MOS-LQO (P.862.1) 8 kHz / 16 kHz Wideband MOS-LQO (P.862.2) 16 kHz without Corrigendum 2 Wideband MOS-LQO (P.862.2) 16 kHz with Corrigendum 2 Accepts stereo inputs ITU reference code (ANSI C) (RMSE = 0.62) (RMSE = 0.61) (Comparison point) (RMSE = 0.56) (Intended?) ludlows/PESQ (Python) (RMSE = 0.03) audiolabs/torch-pesq (PyTorch) (RMSE = 0.15) vBaiCai/python-pesq (Python) Loizou (MATLAB) audiolabs/PESQ (Python)
@@
Figure 1: Mapping from narrowband raw PESQ (P.862) to narrowband MOS-LQO (P.862.1) in the left-hand figure and difference (diff) plot on the right.
@@
The third version of PESQ is P.862.2 [4]. This considers a larger frequency range (50-7000 Hz, named wideband or headphone listening) and is available only for signals with a sampling frequency of 16 kHz. This version only outputs MOS-LQO scores and no raw PESQ scores. Compared to P.862.1, P.862.2 introduces new input filters and a new output mapping to MOS-LQO.
@@
More than a decade after the first edition of P.862.2, the filter coefficients for wideband assessment were acknowledged to be incorrect in P.862 Corrigendum 2 [5]. The corrigendum describes a systematic under-prediction of subjective scores. The under-prediction of P.862.2 is estimated to be $0.8$ on average, and the corrected filter coefficients are provided. Despite the acknowledged errors, P.862.2 without Corrigendum 2 seems to have established itself as the de facto standard in the speech processing community. This could be explained by the lack of an openly available implementation that directly includes Corrigendum 2. We fill this gap by introducing audioalabs/PESQ, as explained in Sec. 3.2.
@@ 3.2 Python and PyTorch
ludlows/PESQ [^2] https://github.com/ludlows/PESQ wraps around the ITU reference code and provides access to P.862.1 and P.862.2 for Python users. It is released as the PyPI package pesq.
audiolabs/PESQ [^3] https://github.com/audiolabs/PESQ is the repository we created by forking from ludlows/PESQ and adding the corrections detailed in Corrigendum 2. It is released as the PyPI package pesqc2. This provides the possibility of easily and reliably computing P.862.2 with Corrigendum 2.
audiolabs/torch-pesq [^4] https://github.com/audiolabs/torch-pesq provides a torch implementation of P.862.2 without Corrigendum 2 that can be readily used as loss function for training DNNs.
@@ 3.3 MATLAB
4 Comparative Analysis
This section quantifies the differences between the different versions (Sec. 4.2) and the different implementations (Sec. 4.3) of PESQ. The audio signals used in these comparisons are detailed in Sec. 4.1.
4.1 Test audio signals
As test audio signals, we use the Open Dataset of Audio Quality (ODAQ) [17]. It is an openly available dataset of stereo audio signals with $44.1$ or $48$ kHz sampling frequency. The dataset was designed to span a heterogeneous range of processing types and quality levels, i.e., from bad to excellent perceived quality. This is particularly useful for our analysis, as it allows us to analyze PESQ scores that span the entire quality scale. For the evaluation in this section, the stereo signals are passively downmixed (i.e., by averaging the two channels) before computing PESQ. The signals are also downsampled to $16$ kHz using librosa [18] with its default parameters, i.e., soxr_hq.
We consider the listening test part of ODAQ that contains 240 audio samples with an average duration of 12.5 seconds. Of these, 80 samples contain speech panned to the phantom center mixed with stereo music, effects, and environmental sounds. Speech signals are processed with 5 different SE systems, 4 of which are DNN-based SE systems from the literature. The SE systems provide clearly distinct quality levels. The non-speech signals mainly include music and are processed by methods simulating quality degradations possibly encountered during audio coding.
Since the goal of this section is to examine how PESQ scores differ for the same input signals across versions and implementations, both speech and non-speech signals are considered. This ensures that PESQ mechanics are evaluated across a broad range of conditions and operating points, including conditions outside of the PESQ application domain. When computing PESQ, we use the reference of the listening test in ODAQ as reference signal for PESQ.
ODAQ also provides ground-truth subjective scores of perceived quality. These are not used in this section. In contrast, they are used in Sec. 5.3, but in that case only speech signals are considered.
4.2 Comparing versions
In this section, we evaluate the differences between PESQ versions based on their implementation in the ITU reference code. We report Pearson’s linear correlation $\rho$, Root Mean Squared Error (RMSE), mean difference, and max absolute difference. RMSEs are summarized in the first row of Table 2.
Figure 2 compares wideband P.862.2 without Corrigendum 2 to narrowband raw PESQ (P.862), using a sampling frequency of 16 kHz. The max difference is $1.82$ and RMSE =0.62=0.62 on the considered data. The trend is similar (hence, the plot is not included) when comparing P.862.2 and P.862.1 (16 kHz), resulting in $\rho=0.89$, RMSE =0.61=0.61, mean difference =0.41=0.41, and max difference =0.92=0.92.
Figure 3 shows the differences between P.862.2 with and without Corrigendum 2. Although the correlation is very strong, the differences in absolute terms are significant with an RMSE =0.56=0.56 and a maximum difference of $1.30$ for the audio signals considered.
Figure 3: PESQ P.862.2 with and without Corrigendum 2.
4.3 Comparing implementations
In this paragraph, we focus on P.862.2 (without Corrigendum 2) and compare different implementations. RMSEs are summarized in the third column in Table 2. Figure 4 compares P.862.2 as output by the ITU reference code and by ludlows/PESQ, showing that they are almost identical. Figure 5 compares audiolabs/torch-pesq and the ITU reference implementation. As acknowledged by the authors of torch-pesq, some items can deviate significantly from the ITU reference implementation. This is because some processing steps were approximated to allow for a differentiable implementation of PESQ. Still, the average error is rather small with RMSE =0.15=0.15 on the data considered.
Figure 4: ludlows/PESQ compared with the ITU reference implementation. Differences are negligible.
Furthermore, our audiolabs/PESQ was compared against the ITU reference implementation with the addition of Corrigendum 2. The differences are negligible (hence, plot not included), with $\rho=1.00$, RMSE =0.01=0.01, mean difference =0.00=0.00, and max absolute difference =0.12=0.12. This confirms that audiolabs/PESQ is a valid out-of-the-box solution for computing P.862.2 with Corrigendum 2.
Figure 5: audiolabs/torch-pesq compared with the ITU reference implementation.
@@ 5.1 Reference implementation
The results shown so far are obtained considering signals downmixed to mono. Indeed, PESQ was developed for mono inputs, and all considered publicly available implementations accept only mono inputs (or batches thereof), with one important exception: The ITU reference code accepts multi-channel inputs. It is not clear whether this is intentional. In the reference code, the reading routines load all input file samples. In practice, when a multi-channel audio file is given as input, a mono signal is created, in which channels are interleaved. As a consequence, the reference implementation processes the interleaved multi-channel signals as a long single-channel signal.
5.2 Strategies for applying PESQ to stereo
The processing of stereo signals by the ITU reference code is referred to as stereo. More intuitive strategies for applying PESQ to stereo signals include:
@@ 5.2 Strategies for applying PESQ to stereo
    Mono-downmixing the stereo audio signal before computing PESQ (mono dmx).
@@
These stereo strategies are compared in Fig. 6 and Fig. 7, in both cases using the ITU reference implementation with the addition of Corrigendum 2 (c2).
5.3 Comparing stereo strategies
To better understand these strategies, we analyzed the correlation with subjective scores of perceived audio quality. As the ground truth, we used the scores collected from two MUSHRA [19] listening tests involving stereo signals with speech panned to the phantom center of the stereo scene:
@@ 5.3 Comparing stereo strategies
    The Dialogue Enhancement (DE) subset of ODAQ containing only speech signals. A more detailed analysis for the correlation on this dataset is given in  [20].
@@
    The results from the listening test described in [21], which compares SE methods applied to real-world broadcast speech signals.
We observed sensible PESQ scores in all cases, with Pearson’s linear correlation coefficients $\geq 0.73$ and reaching $0.85$ in the best cases. However, inconclusive results are observed, and no stereo strategy consistently outperforms the others on both considered datasets. This prevents drawing a definitive conclusion on the overall best strategy.
Figure 6: Comparing strategies for using PESQ (P.862.2 c2) with stereo signals: mono dmx and avg scores.
Figure 7: Stereo signals input to the ITU reference implementation (with c2) against mono-downmixed inputs.
@@ 6 Conclusion
Although replaced by POLQA, withdrawn by the ITU, and outperformed by other quality measures, PESQ remains a widely used speech quality metric. The term PESQ encompasses different versions: narrowband raw PESQ score (P.862), narrowband MOS-LQO (P.862.1), and wideband MOS-LQO (P.862.2), with and without Corrigendum 2.
This paper provides practical guidance on the different versions and implementations of PESQ. Differences between PESQ versions are shown to be particularly significant. Hence, it is recommended to specify the PESQ version and implementation when reporting PESQ scores, as this ensures clarity and reproducibility across studies.
In particular, Corrigendum 2 introduced significant adjustments to P.862.2, but it was not included in any of the publicly available implementations of PESQ. Probably due to this lack, the community seemed to ignore the corrigendum in practice, and P.862.2 without Corrigendum 2 is used most frequently. Our audiolabs/PESQ fills this gap and provides easy and reliable access to P.862.2 with Corrigendum 2, i.e., the last version of PESQ.
Finally, attention is drawn to the likely unintentional behavior of the ITU reference implementation when dealing with multi-channel signals. Alternative strategies are discussed, providing plausible quality scores for stereo signals with speech panned to the center.